An employer interviews 8 people for 6 computer programming positions in a company, and 3 of the 8 people are women. If all 8 are qualified, in how many ways can the employer fill the computer programming positions if exactly 2 are women?
Answer by stanbon(75887)

An employer interviews 8 people for 6 computer programming positions in a company, and 3 of the 8 people are women. If all 8 are qualified, in how many ways can the employer fill the computer programming positions if exactly 2 are women?
---------------------------
# of ways to succeed: 3C2*5C4 = 3*5 = 15
---
# of possible grps of 6:: 8C6 = 8C2 = 28
-----
P(2women and 4men) = 15/28
===================
Cheers,
Stan H.
===================